A Child weighing 25 kg lide down a rope hanging from branch of a tall tree. if the force of friction acting against him is 200 N, the accleration of child is
 
in the FBD of this question. mg acts downward, friction acts upwardly but what about tension ?

Grade:11

1 Answers

Arun
5 years ago
since child of weight is 250N  downward and and friction force 200N act on them.....
 so, net force=250-200=50N

so we known ,

F=ma,

a=f/m

a=50/25

a=2m/s